# Breaking Barriers or Building Dependency? Exploring Team-LLM Collaboration in AI-infused Classroom Debate

### v1

@@ 5.5.2. Cultural Bias in AI-Generated Responses
Participants expressed that AI’s culturally specific examples could lead to understanding challenges and perceptions of cultural bias. P(1) felt that AI provided ”a very bad example” when it referenced cultural contexts that ”people weren’t familiar with,” making these concepts ”relatively abstract.” Additionally, AI would sometimes mention examples such as ”Apple products” or ”Air Jordan shoes”, which P(1) believed reflected cultural bias. He noted, ”Once the examples are grounded in (our) everyday life, our discussions become easier, as people can more easily understand those examples and empathize when facing real-world situations.”
@@ 5.5.3. Formulaic Responses from AI
Our analysis of participant conversations with AI showed that some participants felt AI’s responses were formulaic. P(20) noted, ”AI might respond to a particular argument with three sentences: the first is ’what,’ the second is ’why,’ and the third is a ’summary’, like a classic three-part structure. Every response is like that.” P(1) also stated, ”AI sometimes likes to give obvious and repetitive responses.” Participants noted that AI’s responses to certain questions could feel repetitive, potentially leading to a lack of diversity.
@@ 6. Discussion
In this section, we analyze and discuss the collaboration between teams and AI based on the findings and related works. We also propose design strategies to maximize the benefits while minimizing potential risks. Additionally, we review the limitations of human-AI interfaces in supporting team collaboration and the constraints of existing AI tools in dialogue-based interactions to explore future research directions and potential areas for improvement.
@@ 6.1. Does AI Help or Hinder Learning Objectives in Formal Debate?
In Section 5.2, we explored how AI provides scaffolding across various aspects of formal debate, such as outlining the debate process, crafting rebuttal strategies, and adhering to formal language and etiquette. According to Zone of Proximal Development Theory (Vygotsky, 1978), learners achieve tasks beyond their independent capabilities with appropriate scaffolding. Traditionally, this guidance comes from teachers or self-directed learning through books and online searches. However, AI offers an accessible and dynamic source of scaffolding. Through AI, learners can master low-order thinking, such as acquiring knowledge on rhetoric and decorum, while simultaneously honing higher-order thinking, such as analyzing debate topics, evaluating collected information, and responding to opponents’ arguments under the constraints of a time-sensitive debate setting.
In Section 4.2, Section 4.3 and Section 5.2, we emphasized how AI can provide large volumes of information quickly, enabling learners to gain a more comprehensive understanding of debate topics. By engaging in Exploratory Talk (Mercer, 2002), which involves critical negotiation, learners can deepen their reasoning skills and construct knowledge collaboratively. AI contributes to this process by providing real-time feedback, such as clarifying points or supplementing evidence, thereby improving discussion quality. This integration of AI supports a deeper understanding of complex issues and strengthens learners’ critical thinking skills.
@@ 6.2. Unique Contribution of AI-infused Classroom Debate
In contrast, AI-infused group decision-making relies more on AI to facilitate divergent thinking and generate diverse creative ideas (Yu-Han and Chun-Ching, 2023). Debates emphasize logically cohesive arguments, linguistic precision, rhetorical finesse, and strategic application of reasoning. Additionally, the intensity of debates makes cognitive resources more precious, rendering participants highly sensitive to information overload. Excessive or irrelevant information can disrupt participants’ focus and hinder their responsiveness in such high-pressure environments. However, other AI-infused collaborations typically do not operate under such urgency, they usually have more time to eliminate useless information (Chiang et al., 2024; Yu-Han and Chun-Ching, 2023; Liu et al., 2024a). These cognitive demands highlight the need to effectively manage and filter AI-generated information in fast-paced scenarios, presenting a critical challenge for real-time AI assistance.
Moreover, Park et al. proposed some challenges of using LLMs in higher education, and our study has solved some of them, including ”decreased learning and academic integrity” and ”reduced social interaction” (Park and Ahn, 2024). For the former, in the AI-infused debate, AI does not directly act as a debater, in contrast, it acts as a tutor to provide scaffolds, debate rules, and strategy and also acts as an assistant to provide vast of information for debaters, thus debaters not only avoid reducing learning and academic integrity but can also use tutoring to quickly grasp debate knowledge and frameworks. This approach enables them to efficiently formulate their arguments in high-paced scenarios. For the later challenge, we intentionally stipulate that only one laptop can be accessed and use ChatGPT 3.5 in a group in our study, our findings show that because only a few people can use ChatPT, they developed a clear division of labor and communicate deeply face-to-face or via social media, which promote their social interaction skills. These contributions demonstrate the success of our designed AI-infused debate.
@@ 6.3. Limitations of Current Human-AI Tools in Dialogue-Based Group Collaboration
Design Implication 1: Improving Information Sharing and Communication Flow Addressing the challenges of high-paced and high-intensity collaborative learning scenarios, future team-used AI designs should prioritize developing more intuitive, adaptive interfaces. These interfaces could incorporate multimodal and multisensory technologies to enhance interpersonal communication and facilitate multi-user real-time interaction and information sharing. This design implication aligns with Intersubjectivity (Rogoff, 1990) and Sociocultural Theory (Mercer, 2002).
Our study also found that AI has a short memory for text, resulting in discontinuous conversations and frequent omissions, especially when handling complex or multi-step instructions requiring repeated inputs. Furthermore, AI often lacks a clear stance when dealing with ambiguous viewpoints. When providing examples, AI responses always mix useful information with irrelevant details, which increases the user’s workload of filtering content (Soman et al., 2023). The long-term memory and context-alignment of LLMs are inherently determined by model performance or the framework that LLMs use, and numerous AI scientists are actively working to address this challenge (Wang et al., 2024; Liu et al., 2023a).
Design Implication 2: Enhancing AI’s Output Using Multi-user Selected Information Allow team members to mark and save AI responses they find important or satisfactory during interactions, introducing a feedback mechanism with options such as ”Support,” ”Oppose,” and ”Neutral.” Add interface options for users to specify the stance the AI should adopt, enabling the AI to generate responses that better align with user expectations. This approach allows the AI to reference saved information in subsequent conversations, maintaining continuity in dialogue. Establish a team-level memory repository to consolidate important responses saved by all members, ensuring that all key information is considered during team interactions. Additionally, it provides a convenient search and reference function in the interface, allowing team members and the AI to quickly locate and cite previously saved content, reducing information omissions and minimizing the need for repetitive input.
@@
According to our findings, many participants indicated that AI tools allowed them to quickly obtain large amounts of information, greatly facilitating the retrieval of relevant data when needed. However, the speed and volume of information provided by AI also come with potential side effects, such as the risk of information overload. Participants reported spending a significant amount of time trying to discern valuable information, and some resorted to ”giving up” by completely trusting AI’s responses. This contradicts one of the key characteristics of human-AI collaborative learning: ”mutual understanding” (Huang et al., 2019). At the same time, participants acknowledged that AI’s ability to provide large amounts of content quickly offered beginners an overview and helped them build a preliminary understanding. This highlights AI’s dual nature in terms of information quality.
Design Implication 3: Mitigating Information Overload Providing contextually appropriate explanations depending on the stage of interaction would allow users to extract more relevant insights and promote deeper collaboration. By gradually breaking down complex queries into sub-questions, AI systems can encourage more effective learning and mutual understanding, following the principles of ”mutual understanding” and ”mutual growth” in human-AI interactions (Huang et al., 2019).
@@ 6.5. Moderate AI Dependency in Group Collaboration
Through analyzing participant feedback, this study found that the convenience and rapid information-providing capabilities of AI tools could lead to over-reliance in some cases, reducing participants’ independent thinking. In time-constrained situations, some participants chose to think independently before turning to AI, reflecting their awareness of the dangers of over-reliance on AI tools and a desire to maintain their capacity for independent thought. We also noted that participants’ ”self-awareness” behaviour is related to their confidence levels, which echoes Chong et al.’s research, demonstrating that an individual’s confidence level influences their decision to accept or reject AI suggestions (Chong et al., 2022). Specifically, individuals with higher confidence in their judgment are more likely to question or ignore AI recommendations, while those with lower confidence may over-rely on incorrect AI suggestions. This phenomenon is referred to as ”human self-confidence calibration” (Ma et al., 2024b), which suggests that researchers should assess and calibrate users’ confidence levels before they start using AI tools.
Design Implication 4: Dynamic Adjustment of AI’s Role in Team Collaboration Future AI design should incorporate mechanisms to encourage critical thinking, particularly by adjusting AI engagement based on users’ confidence levels. For highly confident users, the system can pose challenging questions or case analyses to stimulate critical thought, while for less confident users, a gradual approach with guidance and affirmation can help improve their decision-making abilities. This approach fosters a balance between AI assistance and independent thinking. In addition, dynamic AI adjustment features can be integrated into team collaboration, allowing tasks and roles to be flexibly assigned based on each team member’s performance and needs. This method not only balances individual participation in group efforts but also enhances a sense of achievement in collaborative learning. Adjustments can be based on members’ confidence, interaction frequency, contributions, and feedback, ensuring everyone actively engages in tasks aligned with their abilities and interests, ultimately deriving fulfillment from the collaborative process.
@@ 6.6. Limitations and Future Work
This research has revealed several limitations during the study process. Firstly, due to a limited sample size, our results may not fully represent a wide range of collaborative learning scenarios. Additionally, technological maturity has not yet fully met educational needs. The dialogue systems utilized in this study have not been widely implemented in practical study environments, and both the technical maturity and user adaptability remain to be verified. This limitation might impact the practical applicability of our findings. Data security in human-AI interaction is another significant concern. Existing studies have identified biases and stereotypes in the outputs of Large Language Models (LLMs) (Bai et al., 2024; Gupta et al., 2023). Furthermore, our study lacks diversity since our participants were primarily from specific age and academic groups. The classroom types explored may also limit the applicability of our findings across broader educational formats. The absence of long-term tracking of technology interventions makes it challenging to assess their enduring impact, and the potential novelty effect introduced by technologies like ChatGPT must also be addressed in future work (Kim et al., 2024). During the classroom debate experiments, ChatGPT-4 had just been released and was relatively slow, making it unsuitable for high-paced debate settings. Faster-response large language models (LLMs) like GPT-4o had not yet been released. Thus, ChatGPT-3.5 was chosen for its faster response times. While the focus of the study was not on the performance of LLMs but rather on the interaction patterns within AI-infused team collaboration, it is undeniable that the capabilities of LLMs could influence the information-gathering process and the degree of response hallucination (Huang et al., 2023), ultimately impacting the outcomes of the debates. Future research could build upon this work by employing state-of-the-art LLMs to explore these dynamics further. Additionally, all participants are Chinese, which means they operate within a specific cultural context when gathering information on platforms like Zhihu and Xiaohongshu. This cultural specificity similarly influences their use of AI. Consequently, addressing how different cultural norms and practices in information retrieval can be incorporated into AI-infused collaborative learning represents both a limitation of this study and a significant avenue for future research. In the future, a broader range of cultural norms need to be included in the research, thereby improving the generalizability of the findings. Finally, although classroom debate activities were already part of the course’s regular teaching practices, with occasional invitations extended to external teachers or students as observers, the setup in this study does not differ significantly from the usual classroom environment. However, we cannot rule out the potential influence of the Hawthorne effect (Adair, 1984), which is a common challenge in observational research.
@@ 7. Conclusion
This study investigates interactive behaviours in human-AI hybrid collaborative learning through the creation of five-to-five classroom debates involving 22 students. The findings indicate that various collaborative modes between team members and AI emerged during the interaction. AI’s involvement altered how humans posed questions and utilized AI-generated answers, fostering a division of labour within teams. Additionally, different methods of using tools for recording and exchanging views were observed. These processes helped groups build consensus, filter information, deepen understanding, and form diverse perspectives, thereby enhancing team collaboration. Moreover, AI significantly reduced users’ psychological burden and communication barriers within groups, lowered the entry threshold for novice debaters—such as addressing cognitive biases—stimulating individual thinking, and sparked discussions between participants.
However, the study also reveals several challenges in the process of human-AI collaborative learning. AI may lead to information overload, induce dependency, limit human autonomy, and cause the perceived low quality of AI-generated content, which in turn leads to cognitive inertia, weakens personal initiative, and restricts the generation of new ideas. To maximize the benefits of AI while mitigating associated risks, the study proposes design strategies, analyzes the limitations of current human-computer interfaces in collaborative learning, and discusses the text memory constraints of existing AI tools in dialogue tasks. Furthermore, it highlights the impact of AI-driven task allocation on participants’ motivation and engagement in collaborative learning. The study’s limitations include a small sample size, insufficient technological maturity, data security concerns, the homogeneity of the user group, and a lack of long-term impact assessment. These factors provide multiple dimensions for the expansion and deepening of future research.

### v2

@@ 5.5.2. Cultural Bias in AI-Generated Responses
Participants expressed that AI’s culturally specific examples could lead to confusion and perceptions of cultural bias. P(1) felt that AI provided ”a very bad example” when it referenced cultural contexts that ”people weren’t familiar with,” making these concepts ”relatively abstract.” Additionally, AI would sometimes mention examples such as ”Apple products” or ”Air Jordan shoes”, which P(1) believed reflected cultural bias. He noted, ”Once the examples are grounded in (our) everyday life, our discussions become easier, as people can more easily understand those examples and empathize when facing real-world situations.”
@@ 5.5.3. Formulaic Responses from AI
Our analysis of participant conversations with AI showed that some participants felt AI’s responses were formulaic. P(20) noted, ”AI might respond to a particular argument with three sentences: the first is ’what,’ the second is ’why,’ and the third is a ’summary,’ like a classic three-part structure. Every response is like that.” P(1) also stated, ”AI sometimes likes to give obvious and repetitive responses.” Participants noted that AI’s responses to certain questions could feel rigid and mechanical, potentially leading to a lack of diversity.
@@ 6. DISCUSSION
In this section, we analyze and discuss the collaboration between teams and AI based on the findings and related works. We also propose several design strategies to maximize the benefits while minimizing potential risks. Additionally, we review the limitations of human-AI interfaces in supporting team collaboration and the constraints of existing AI tools in text-based interactions to explore future research directions and potential areas for improvement.
@@ 6.1. Does AI Help or Hinder Learning Objectives in Formal Debate?
In Section 5.2, we explored how AI provides scaffolding across various aspects of formal debate, such as outlining the debate process, crafting rebuttal strategies, and adhering to formal language and etiquette. According to Zone of Proximal Development Theory (Vygotsky, 1978), learners achieve tasks beyond their independent capabilities through collaboration and guidance in a socially constructed learning space. This guidance traditionally comes from teachers or peers, who act as a more knowledgeable other (MKO), or from self-directed learning through books and online searches. In this study, AI plays an unconventional yet intriguing role as an MKO. While it would not be accurate to claim that AI is more dynamic than a human collaborator, the flexibility and vast knowledge base of AI make it a unique tool for scaffolding in diverse contexts. This study shows AI’s potential to create a dynamic and inclusive learning environment, extending opportunities to learners regardless of their physical or social access to traditional resources.
In Section 4.2, Section 4.3, and Section 5.2, we emphasized how AI can provide large volumes of information quickly, enabling learners to gain a more comprehensive understanding of debate topics. Learners can deepen their reasoning skills and construct knowledge collaboratively by engaging in Exploratory Talk (Mercer, 2002), which involves critical negotiation. AI contributes to this process by providing real-time feedback, such as clarifying points or supplementing evidence, thereby improving discussion quality. This integration of AI supports a deeper understanding of complex issues and strengthens learners’ critical thinking skills.
@@ 6.2. Unique Contribution of AI-infused Classroom Debate
In contrast, AI-infused group decision-making relies more on AI to facilitate divergent thinking and generate diverse creative ideas (Yu-Han and Chun-Ching, 2023). Debates emphasize logically cohesive arguments, linguistic precision, rhetorical finesse, and strategic application of reasoning. Additionally, the intensity of debates makes cognitive resources more precious, rendering participants highly sensitive to information overload. Excessive or irrelevant information can disrupt participants’ focus and hinder their responsiveness in high-pressure environments. However, other AI-infused collaborations typically do not operate under such urgency, and they usually have more time to eliminate useless information (Chiang et al., 2024; Yu-Han and Chun-Ching, 2023; Liu et al., 2024a). These cognitive demands highlight the need to effectively manage and filter AI-generated information in fast-paced scenarios, presenting a critical challenge for real-time AI assistance.
Moreover, Park et al. proposed some challenges of using LLMs in higher education, and our study has solved some of them, including ”decreased learning and academic integrity” and ”reduced social interaction” (Park and Ahn, 2024). For the former, in the AI-infused debate, AI does not directly act as a debater. In contrast, it acts as a tutor to provide scaffolds, debate rules, and strategy and as an assistant to provide vast information for debaters. Thus, debaters avoid reducing learning and academic integrity and can use tutoring to quickly grasp debate knowledge and frameworks. This approach enables them to efficiently formulate their arguments in high-paced scenarios. For the later challenge, we intentionally stipulate that only one laptop can be accessed and use ChatGPT 3.5 in a group in our study. Our findings show that because only a few people can use ChatGPT, they developed a clear division of labor and communicate deeply face-to-face or via social media, which promotes their social interaction skills. These contributions demonstrate the success of our designed AI-infused debate.
@@ 6.3. Limitations of Current Human-AI Tools in Dialogue-Based Group Collaboration
Design Implication 1: Improving Information Sharing and Communication Flow
To address the challenges of high-paced and high-intensity collaborative learning, future LLM-based systems should move beyond isolated one-on-one interactions and manual workflows like copying outputs into shared online documents. Instead, they should offer integrated, adaptive interfaces for seamless team collaboration. These interfaces could support multimodal inputs, enabling users to communicate naturally through text, voice, images, and other forms within the same platform.
Real-time interaction and shared workspaces should replace fragmented processes, allowing team members to collaboratively co-create, discuss, and edit AI-generated outputs. By incorporating features like synchronized updates, role-based views, and tools for annotating or summarizing multimodal content, such systems can streamline communication, enhance interpersonal connections, and meet the demands of fast-paced team environments. These designs align with Intersubjectivity (Rogoff, 1990) and Sociocultural Theory (Mercer, 2002).
Our study also found that AI has a limited memory for text, resulting in discontinuous conversations and frequent omissions, especially when handling complex or multi-step instructions requiring repeated inputs. Furthermore, AI often lacks a clear stance when dealing with ambiguous viewpoints. When providing examples, AI responses always mix useful information with irrelevant details, which increases the user’s workload of filtering content (Soman et al., 2023). Although the long-term memory and context-alignment of LLMs are inherently determined by the model training process or the framework that LLMs use, and numerous AI scientists are actively working to address this challenge (Wang et al., 2024; Liu et al., 2023a), we can adopt specific strategies to make their responses more focused and efficient without compromising the LLMs’ performance.
Design Implication 2: Enhancing AI’s Output Using Team-members Selected Information
Allow team members to mark and save AI responses they find important or satisfactory during interactions, introducing a feedback mechanism with options such as ”Support,” ”Oppose,” and ”Neutral.” Add interface options for users to specify the stance the AI should adopt, enabling the AI to generate responses that better align with user expectations. This approach allows the AI to reference saved information in subsequent conversations, maintaining continuity in dialogue. Establish a team-level memory repository to consolidate meaningful responses saved by all members, ensuring that all key information is considered during team interactions.
Additionally, it should provide convenient search and reference functions in the interface, allowing team members and the AI to quickly locate and cite previously saved content, reducing information omissions and minimizing the need for repetitive input.
@@
According to our findings, many participants indicated that AI tools allowed them to quickly obtain large amounts of information, greatly facilitating the retrieval of relevant data when needed. However, the volume of information AI provides also comes with the potential risk of information overload. Participants reported spending significant time trying to discern valuable information, and some resorted to ”giving up” by completely trusting AI’s responses. This contradicts one of the key characteristics of human-AI collaborative learning: ”mutual understanding” (Huang et al., 2019). At the same time, participants acknowledged that AI’s ability to provide large amounts of content quickly offered beginners an overview and helped them build a preliminary understanding. This highlights AI’s dual nature in terms of information quality.
Design Implication 3: Mitigating Information Overload
Providing contextually appropriate explanations based on the stage of interaction allows users to extract more relevant insights and fosters deeper collaboration. For example, before responding, a system could offer preset options such as ”Concise,” ”Moderate,” or ”Detailed.” Users could then select the desired level of detail according to their needs, avoiding situations where they must wait for the AI to return an excessive amount of information and subsequently summarize it themselves or request the AI to summarize it again.
When addressing complex problems, AI should be designed to break down the issue into several sub-questions, progressing from simple to complex. This step-by-step approach would guide users in thinking through more intricate aspects of the problem while avoiding the overwhelming presentation of excessive information all at once. Such a design ensures that users can better grasp the key points of the response without being distracted by an overload of details, following the principles of ”mutual understanding” and ”mutual growth” in human-AI interactions (Huang et al., 2019).
@@ 6.5. Moderate AI Dependency in Group Collaboration
Through analyzing participant feedback, this study found that the convenience and rapid information-providing capabilities of AI tools could sometimes lead to over-reliance, reducing participants’ independent thinking. In time-constrained situations, some participants chose to think independently before turning to AI, reflecting their awareness of the dangers of over-reliance on AI tools and a desire to maintain their capacity for independent thought. We also noted that participants’ ”self-awareness” behavior is related to their confidence levels, which echoes Chong et al.’s research, demonstrating that an individual’s confidence level influences their decision to accept or reject AI suggestions (Chong et al., 2022). Specifically, individuals with higher confidence in their judgment are more likely to question or ignore AI recommendations, while those with lower confidence may over-rely on incorrect AI suggestions. This phenomenon is referred to as ”human self-confidence calibration” (Ma et al., 2024b), which suggests that researchers should assess and calibrate users’ confidence levels before they start using AI tools.
Design Implication 4: Dynamic Adjustment of AI’s Role in Team Collaboration
Future AI design should incorporate mechanisms to encourage critical thinking, particularly by adjusting AI engagement based on users’ confidence levels. For highly confident users, the system can pose challenging questions or case analyses to stimulate critical thought, while for less confident users, a gradual approach with guidance and encouragement can help improve their decision-making and problem-solving abilities. This approach fosters a balance between AI assistance and independent thinking.
In addition, dynamic AI adjustment features can be integrated into team collaboration, allowing tasks and roles to be flexibly assigned based on each team member’s performance and needs. Adjustments can be based on members’ confidence, interaction frequency, contributions, and feedback, ensuring everyone actively engages in tasks aligned with their abilities and interests, ultimately deriving fulfillment from the collaborative process. This method balances individual participation in group efforts and enhances a sense of achievement in collaborative learning.
@@ 6.6. Limitations and Future Work
This research has revealed several limitations during the study process. Firstly, our results may not fully represent a wide range of collaborative learning scenarios due to a limited sample size. Additionally, technological maturity has not yet fully met educational needs. The dialogue systems utilized in this study have not been widely implemented in practical study environments, and both the technical maturity and user adaptability remain to be verified. This limitation might impact the practical applicability of our findings. Data security in human-AI interaction is another significant concern. Existing studies have identified biases and stereotypes in the outputs of Large Language Models (LLMs) (Bai et al., 2024; Gupta et al., 2023). Furthermore, our study lacks diversity since our participants were primarily from specific age and academic groups. The classroom types explored may also limit the applicability of our findings across broader educational formats. The absence of long-term tracking of technology interventions makes it challenging to assess their enduring impact, and the potential novelty effect introduced by technologies like ChatGPT must also be addressed in future work (Kim et al., 2024). ChatGPT-4 had just been released during the classroom debate experiments and was relatively slow, making it unsuitable for high-paced debate settings. Faster-response large language models (LLMs) like GPT-4o had not yet been released. Thus, ChatGPT-3.5 was chosen for its faster response times. While the focus of the study was not on the performance of LLMs but rather on the interaction patterns within AI-infused team collaboration, it is undeniable that the capabilities of LLMs could influence the information-gathering process and the degree of response hallucination (Huang et al., 2023), ultimately impacting the outcomes of the debates. Future research could build upon this work by employing state-of-the-art LLMs to explore these dynamics further. Additionally, all participants are Chinese, which means they operate within a specific cultural context when gathering information on platforms like Zhihu and RedNote. This cultural specificity similarly influences their use of AI. Consequently, addressing how different cultural norms and practices in information retrieval can be incorporated into AI-infused collaborative learning represents a limitation of this study and a significant avenue for future research. In the future, a broader range of cultural norms needs to be included in the research, thereby improving the generalizability of the findings. Finally, although classroom debate activities were already part of the course’s regular teaching practices, with occasional invitations extended to external teachers or students as observers, the setup in this study does not differ significantly from the usual classroom environment. However, we cannot rule out the potential influence of the Hawthorne effect (Adair, 1984), a common challenge in observational research.
@@ 7. CONCLUSION
This study investigates interactive behaviors in human-AI hybrid collaborative learning by creating five-to-five classroom debates involving 22 students. The findings indicate various collaborative modes emerged between team members and AI during the interaction. AI’s involvement altered how humans posed questions and utilized AI-generated answers, fostering a division of labor within teams. Additionally, different methods of using tools for recording and exchanging views were observed. These processes helped groups build consensus, filter information, deepen understanding, and form diverse perspectives, enhancing team collaboration. Moreover, AI significantly reduced users’ psychological burden and communication barriers within groups, lowered the entry threshold for novice debaters—such as addressing cognitive biases—stimulating individual thinking, and sparked discussions between participants. However, the study also reveals several challenges in human-AI collaborative learning. AI may lead to information overload, induce dependency, limit human autonomy, and cause the perceived low quality of AI-generated content. This leads to cognitive inertia, weakens personal initiative, and restricts the generation of new ideas. To maximize the benefits of AI while mitigating associated risks, the study proposes design strategies, analyzes the limitations of current human-computer interfaces in collaborative learning, and discusses the text memory constraints of existing AI tools in dialogue tasks. Furthermore, it highlights the impact of AI-driven task allocation on participants’ motivation and engagement in collaborative learning. The study’s limitations include a small sample size, insufficient technological maturity, data security concerns, the homogeneity of the user group, and a lack of long-term impact assessment. These factors provide multiple dimensions for the expansion and deepening of future research.
ACKNOWLEDGMENTS
We sincerely appreciate the involvement of all the volunteers who participated in this study and the 22 students who contributed to the project experiments for their efforts and support in data collection and conducting interviews throughout the research. We also greatly thank the reviewers for their invaluable feedback and suggestions. This work is supported by the National Natural Science Foundation of China (NSFC) Youth Grant (Grant No. 62307024).